OpenStudy (anonymous):

Find the derivative of the integral from x to x^4.

OpenStudy (anonymous):

\[f(x)=\int\limits_{x}^{x^4}t^3dt\]

OpenStudy (anonymous):

The Second Fundamental Theorem of Calculus states: if \[ f(x)=\int\limits_{a}^{x}g(t)dt\] then \[f'(x)= g(x)\]

OpenStudy (anonymous):

I separated the integral : \[\huge f(x)= -\int\limits_{a}^{x}t^3dt+\int\limits_{a}^{x^4}t^3dt\] But now I'm stuck.

OpenStudy (anonymous):

\[\int\limits\limits_{x}^{x^4}t^3dt=\left[ t^4/4 \right]= {(x^4)^4-x^4}/4\]

OpenStudy (anonymous):

That isn't explicitly using the Second Fundamental Theorem though. And it isn't correct.

OpenStudy (anonymous):

y,

OpenStudy (anonymous):

you aren't accounting for the fact that the bottom boundary is x and not a.

OpenStudy (anonymous):

a is constant whereas x isn't.

OpenStudy (anonymous):

but wat difference does it make, if u integrate where u stuck , u will ge the same expression

OpenStudy (anonymous):

The integral will be different from 2-16 than it will be from 3-81.

OpenStudy (anonymous):

f'(x) = -t^3 +4x^3t^3

OpenStudy (anonymous):

Notice your variables don't match...?

OpenStudy (anonymous):

oh, t should be x

OpenStudy (anonymous):

How did you get that answer? It doesn't make sense

OpenStudy (anonymous):

Can you use the equation tool? I cant read that.

OpenStudy (anonymous):

sr, I dont know that tool =.=

OpenStudy (anonymous):

It's right beside the draw tool T.T

OpenStudy (anonymous):

just substitute the endpoint of x with t and reverse. then use the theorem to get answer

OpenStudy (anonymous):

ps: I'm a member for 1 day, so, sorry about that

OpenStudy (anonymous):

I am too. Just a minute let me try that.

OpenStudy (anonymous):

No. That isn't right.

OpenStudy (anonymous):

what you mean, the final answer is -x^3+4x^6 ?

OpenStudy (anonymous):

No, it isnt.

OpenStudy (anonymous):

I'll just ask my prof tomorrow if nobody can help me.

OpenStudy (anonymous):

i referred to a formula, it's called Leibniz's rule